OpenStudy (anonymous):

A biochemist is testing two types of drugs that induce cell growth and replication. She starts with 1000 cells for each of the two drugs tested. Drug A raises the number of cells by 200 cells each day. Drug B induces an 8% increase in the number of cells each day. Which statement is true about the number of cells after 30 days?

OpenStudy (anonymous):

200*30=6000+1000 (original cells)=7000 1000*1.8^30=what? use a calculator.

OpenStudy (anonymous):

Whoa, 1.08^30, I hope I caught that in time

OpenStudy (anonymous):

Okay, that will equal 10062.7, or 10063 in the case of percentage growth